OpenStudy (anonymous):

Harris has a triangular flag. He wants to cut out a similar flag that has dimensions one-fifth of the original flag. If the area of the original flag is 1250 square inches, what is the area of the new flag that Harris cuts out?

OpenStudy (agreene):

one-fifth or 1/5 = 0.2 in decimal form; thus: 1250*0.2 = 250 in^2 is the area of the smaller flag.

OpenStudy (anonymous):

These were my choices 10 in2 25 in2 50 in2 65 in2

OpenStudy (anonymous):

So, 25?

OpenStudy (anonymous):

no, they want the new area with each dimension reduced to 1/5 the original, the area does not reduce to 1/5 however

OpenStudy (phi):

Qwerty, why don't you think about it?  all the sides get smaller by 1/5

OpenStudy (phi):

so what is the area of the smaller triangle?

OpenStudy (agreene):

yeah, i misread the question. the dimensions, not the area reduce.

Directrix (directrix):

50 in^2 --> I would go with this answer because the square of the scale factor of two similar polygons is the ratio of their areas. (1/5) ^2 = A/1250 1/25 = A/ 1250 25A = 1250 A = 50

OpenStudy (anonymous):

Yes I see that now, I tried to smaller the dimensions but I had to figure out what he original dimensiions were

OpenStudy (anonymous):

you don't need the original dimensions

OpenStudy (anonymous):

So, if my areaa is 1250 I have to figure out what 1/5 of the dimensions are so tat they equal the 1250?

OpenStudy (anonymous):